# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1

### v1

@@ Abstract
In this work, we bridge the gap between wearable sensor technology and personalized AI assistants by enabling Large Language Models (LLMs) to understand time-series tasks like human activity recognition (HAR). Despite the strong reasoning and generalization capabilities of LLMs, leveraging them for sensor data tasks remains largely unexplored. This gap stems from challenges like the lack of semantic context in time-series data, computational limitations, and LLMs’ difficulty processing numerical inputs. To address these issues, we introduce SensorLLM, a two-stage framework to unlock LLMs’ potential for sensor data tasks. In the Sensor-Language Alignment Stage, we introduce special tokens for each sensor channel and automatically generate trend-descriptive text to align sensor data with textual inputs, enabling SensorLLM to capture numerical changes, channel-specific information, and sensor data of varying lengths—capabilities that existing LLMs typically struggle with, all without the need for human annotations. Next, in Task-Aware Tuning Stage, we refine the model for HAR classification using the frozen LLM and alignment module, achieving performance on par with or surpassing state-of-the-art models. We further demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, enabling it to generalize across diverse datasets for HAR tasks. We strongly believe our work lays the stepstone for future time-series and text alignment research, offering a path toward foundation models for sensor data. Our codes are available at https://github.com/zechenli03/SensorLLM.
@@ 1 Introduction
Human Activity Recognition (HAR) is a time-series (TS) classification task that involves mapping sensor data, such as accelerometer and gyroscope signals, to human activities. Models such as LSTM (Guan & Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez & Roggen, 2016) have gained popularity in HAR due to their ability to learn high-level features. However, these models are typically task-specific and struggle to scale across varying sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Wu et al., 2023b; Yin et al., 2023), including text and images. In this work, we explore how LLMs can unlock new possibilities for HAR by improving the interpretation of sensor data and offering a more scalable, flexible solution to this challenging task.
To enable LLMs to understand effectively and process sensor data (Jin et al., 2023), there are two primary approaches. The first approach involves pretraining or fine-tuning LLMs on sensor data (Zhou et al., 2023b), optimizing them to handle time-series inputs directly. However, this method presents significant challenges, particularly related to computational resource requirements and data limitations. Pretraining LLMs on such data demands immense computational resources, and the limited availability of labeled sensor data, especially for rare activities, leads to class imbalance, making the training process difficult and inefficient. The second approach to enabling LLMs to understand sensor data focuses on using zero-shot and few-shot prompting, where sensor data is transformed into textual formats that LLMs can process without retraining (Ji et al., 2024). However, this introduces several challenges. One major issue is the numerical and temporal nature of data, which poses challenges due to how LLMs tokenize numerical inputs (Nate Gruver & Wilson, 2023). LLM tokenizers, designed for text, are not optimized for numerical values. They struggle with recognizing patterns and contextual nuances, treating consecutive values as independent tokens and ignoring temporal correlations. Sensor data, characterized by its continuous and extensive nature, often exceeds the maximum sequence length, resulting in truncation and potential loss of vital information. Another challenge lies in converting sensor data into text-like embeddings while preserving high-level time-series features, as this transformation risks losing the temporal relationships crucial for complex activity recognition (Spathis & Kawsar, 2024). Handling multi-channel sensor data further complicates this process, as encoding multi-dimensional information in a way that LLMs can interpret effectively is difficult cause the input for LLMs is univariate. Moreover, designing effective prompts that enable LLMs to interpret sensor readings, detect trends, and classify activities is complex, especially when dealing with intricate time-series patterns (Liu et al., 2023b).
To address these challenges, we introduce a novel approach for aligning sensor data with natural language, allowing LLMs to interactively analyze sensor data through text-based commands (see Figure 1) without requiring any modifications to the LLM itself. A significant challenge in this context is managing the annotation bottleneck, which is further exacerbated by the inherent heterogeneity and complexity of sensor data. Unlike image-text pairs that can be more straightforwardly aligned, sensor data comprises multiple modalities with diverse characteristics, making the interpretation and alignment process considerably more difficult. Previous approaches (Jin et al., 2024a; Sun et al., 2024b) have often used condensed text prototypes for alignment. While these methods provide a solution, they face limitations in interpretability, as the mapping between numerical time-series data and abstract text prototypes is not always intuitive. Moreover, such methods often require multiple experiments to select appropriate text prototypes, adding complexity to the process.
@@
In this work, we propose a novel approach that automatically generates descriptive text from time-series data, removing the need for manual annotations. These descriptions are generated using statistical analyses or simple descriptions of numerical trends. We employ pre-defined templates to automate the generation of descriptive texts, offering a scalable, precise, and interpretable solution for aligning time-series data with text in our SensorLLM framework. SensorLLM introduces a two-stage framework process that has been proven effective across various domains  (Xu et al., 2024; Huang et al., 2023; Jiang et al., 2024). In Sensor-Language Alignment Stage, we generate automatic question-answer pairs to align sensor data with text, providing a scalable and interpretable solution. To preserve the temporal features of the sensor data, we use a pretrained encoder to generate sensor embeddings, which are then mapped into a space that the LLM can interpret. This avoids the issues caused by using a text-specific embedding model used by LLM. Furthermore, we introduce special tokens for each sensor channel, enabling the LLM to capture and interpret multi-channel data effectively. This ensures that both time-series characteristics and sensor-specific information are fully aligned with the LLM’s interpretative capabilities. In Task-Aware Tuning Stage, the aligned data is leveraged to address downstream tasks HAR, utilizing the LLM’s reasoning abilities while keeping its parameters frozen. This alignment framework enhances the capabilities of LLMs beyond their original pretraining. Our work demonstrates the effectiveness of SensorLLM, addressing the questions raised by Tan et al. (2024) regarding the broader applicability of LLMs for time-series data. To our knowledge, this is the first approach to incorporate sensor data into LLMs for sensor data analysis and HAR tasks. The key contributions of this work are as follows:
  1. 1.
@@
    TS-Language Alignment: We introduce a human-intuitive approach for aligning time-series data with automatically generated text, removing the need for manual annotations. By leveraging text similarity metrics and human evaluations, we demonstrate that SensorLLM effectively captures both time-series patterns and channel-specific information, enabling robust multimodal understanding.
  2. 2.
@@
    Task-aware Sensor-Language Model: SensorLLM achieves competitive performance across four HAR tasks, either surpassing or matching SOTA models. This success is driven by our modality alignment and the inclusion of task-specific prompts, both of which significantly enhance SensorLLM’s ability to interpret and classify sensor data effectively.
  3. 3.
@@
    Cross-Dataset Generalization: We demonstrate that SensorLLM maintains superior performance in Task-Aware Tuning Stage, even when applied to datasets distinct from those used in alignment stage, underscoring the model’s robustness and generalizability in HAR tasks.
@@ 2 Related Work
In this section, we explore recent works that utilize LLMs for human-centric tasks, such as human activity recognition and mobile health sensing. Current approaches can be categorized based on two main aspects: (1) how they integrate numerical time-series data with LLMs, and (2) the specific end task, whether it is time-series forecasting, classification, etc. Zhang et al. (2024) provides a comprehensive review of the existing methodologies that leverage LLM capabilities in processing time-series data. Several works investigated using LLM for numerical data by treating time-series signals as raw text. They used the same tokenization and embedding models as text such as PromptCast  (Xue & Salim, 2023) and LLMTime (Gruver et al., 2024). However, despite LLM’s versatility, they are inherently designed for processing sequential text data and not for heterogeneous time-series numerical data. The challenge lies in enabling LLMs to interpret and learn patterns within multimodal time-series data effectively. As highlighted by Spathis & Kawsar (2024), one significant barrier is the absence of suitable tokenizers for numerical data. LLMs typically employ tokenizers that segment text into smaller units, but these tokenizers are not optimized for representing numerical values. Consequently, they may struggle to recognise repetitive patterns and context, often treating consecutive values as independent tokens and overlooking the crucial temporal relationships inherent in time-series data. To extract more meaningful insights from numerical data before feeding them to LLM, several works proposed the use of time-series encoders and tried to align the embedding space with text embedding models. Liu et al. (2024a) aligns ECG data with corresponding textual descriptions using contrastive learning techniques. Similar approaches have been investigated for aligning motion/IoT sensors with text (Zhou et al., 2023c; Haresamudram et al., 2024; Moon et al., 2023; Xia et al., 2024). Please refer to Appendix A.1 for a more in-depth review of existing works.
@@ 3 Methods
In this work, we propose a novel approach to align sensor data with descriptive text using automatically generated question-answer pairs. The aim of this work is to develop a multimodal foundational model with reasoning capabilities to address challenges in analyzing wearable sensor data. As shown in Figure 2, our SensorLLM comprises three core components: 1) a pretrained LLM, 2) a pretrained TS embedder, and 3) an alignment module MLP. The three core modules cooperate in a two-stage framework. In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text based on user instructions (Liu et al., 2023a) and questions. In Task-Aware Tuning Stage, a classification model is used to perform HAR and maintain an end-to-end alignment. Notably, only the parameters of the MLP in both stages, as well as the classifier in the second stage, are updated during training, while the backbone LLM and TS embedder remain frozen. Consequently, only 5.67% of the parameters in stage 1 and 6% in stage 2 are trainable, ensuring a highly efficient and lightweight training process.
@@ 3.1 Sensor-Text Data Generation
A key challenge in aligning time-series data with text for LLM-based tasks is effectively bridging these two modalities. Unlike other multimodal tasks, time-series data typically lacks rich semantic textual information beyond class labels, making manual annotation impractical  (Deldari et al., 2024; Haresamudram et al., 2024). While prior works have used predefined text prototypes for alignment (Sun et al., 2024a; Jin et al., 2024a), we aim to guide the model toward a more human-intended understanding of the data. We argue that time-series data inherently contains semantic information that can be expressed through descriptive text, from simple numerical trends to more complex insights derived from statistical analysis. Our approach focuses on automatically generating descriptive text based on observed trends and fluctuations in the input data. To achieve this, we use a variety of predefined templates to generate diverse question-answer (QA) pairs that describe trend changes in the sensor data, ensuring both accuracy and scalability, while avoiding the potential errors of manual annotation. These templates are randomly combined to produce diverse QA pairs. For example:
@@
where T and S represent placeholders for a specific trend and sensor type, respectively, and t correspond to numerical values. These automatically generated QA pairs offer a scalable and interpretable method for aligning sensor data and text in the latent space, facilitating more effective understanding and interaction with LLMs. See Appendix A.2 for detailed templates.
@@ 3.2 Sensor-Language Alignment
As shown in Figure 2 (a), Sensor-Language Alignment stage employs a generative model designed to create multi-modal sentences that combine single-channel sensor readings with textual descriptions. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ indicates the number of sensor channels, and $T$ represents the total number of time steps in the data sequence. Each individual channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to maintain the uniqueness of information captured by each channel. The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ represents the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range. The variation in the length of segments allows the model to effectively learn from a variety of temporal patterns and trend variations across different segments. By learning from segments with different lengths, the generative model becomes more adept at handling the diverse characteristics of sensor data, ensuring that both long-term trends and short-term fluctuations are adequately represented in the generated multi-modal sentences. We employ the Chronos (Ansari et al., 2024) as TS embedder, which generates the segment embedding $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and the $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization. Unlike the original Chronos implementation, which uses mean normalization A.4, our approach utilizes instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ for improved performance. We use Llama3-8b (Touvron et al., 2023) as our LLM backbone.
@@
Alignment Module. To transform the time-series embeddings $\hat{x}{s}$ into text embeddings $\hat{a}_{s}\in\mathbb{R}^{(l+1)\times D}$ for further downstream tasks, we introduce an alignment projection module. This module is implemented as a multi-layer perceptron (MLP), which first maps the sensor embeddings to an intermediate latent space $d_{m}$ and then projects them to the final target dimension $D$. Formally, the projection can be written as: $\hat{a}_{s}=\mathbf{W}_{2}\cdot\sigma(\mathbf{W}_{1}\hat{x}_{s}+\mathbf{b}_{1})+\mathbf{b}_{2}$ where $\mathbf{W}1\in\mathbb{R}^{d_{m}\times d_{ts}}$ and $\mathbf{W}2\in\mathbb{R}^{D\times d_{m}}$ are learnable weights, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are biases, and $\sigma$ is GELU activation function (Hendrycks & Gimpel, 2016). This projection ensures that the transformed embeddings $\hat{a}_{s}$ are semantically aligned with the text embedding space, making them suitable for subsequent tasks such as generation or classification.
@@
Input Embedding. To enable the LLM to process sensor channel information and align sensor features with text tokens, we introduce two special tokens for each channel (e.g.,<x_acc_start> and <x_acc_end> for the x-axis accelerometer), extending the LLM’s embedding matrix from $\mathbf{E}\in\mathbb{R}^{V\times D}$ to $\mathbf{E}\in\mathbb{R}^{V^{\prime}\times D}$, where $V^{\prime}=V+2c$ and $V$ is the vocabulary size, and $c$ is the number of channels. Then the special token embeddings are concatenated with the aligned sensor embeddings. Finally, the combined sensor embeddings $\hat{o}_{s}\in\mathbb{R}^{(l+3)\times D}$ are concatenated with instruction and question embeddings to form the final input $\hat{z}\in\mathbb{R}^{k\times D}$, where $k$ is the total number of tokens.
@@
Loss Function. Our SensorLLM processes a mixed sequence of tokens, $\mathbf{Z}_{s}=\{z_{s}^{i}\}_{i=1}^{K}$, consisting of both sensor and text embeddings, and generates an output sequence, $\mathbf{Z}_{t}=\{z_{t}^{i}\}_{i=1}^{N}$, where $z_{s}^{i}$ and $z_{t}^{i}\in V^{\prime}$, and $K$ and $N$ represent the number of input and output tokens, respectively. The generative model is trained using a causal language modeling objective, where the next token in the sequence is predicted based on all preceding tokens. To optimize this, we minimize the negative log-likelihood of the generated text tokens ($\mathcal{L}_{gen}=-\sum_{i=0}^{N-1}\log P(z_{t}^{i}|Z_{t}^{<i},z_{s})$). The loss is computed solely for output tokens, excluding those corresponding to instructions, questions, channel special tokens, and sensor data. This ensures that SensorLLM effectively integrates sensor and text embeddings, enabling the generation of coherent and contextually appropriate responses.

### v2

# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2

@@ Abstract
We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from sensor data. Despite their strong reasoning and generalization capabilities, LLMs remain underutilized for motion sensor data due to the lack of semantic context in time-series, computational constraints, and challenges in processing numerical inputs. SensorLLM addresses these limitations through a Sensor-Language Alignment stage, where we introduce special tokens for each sensor channel and automatically generate textual trend descriptions. This alignment enables LLMs to capture numerical variations, channel-specific features, and data of varying durations—without requiring human annotations. In the subsequent Task-Aware Tuning stage, we refine the model for HAR classification, achieving performance that matches or surpasses state-of-the-art methods. Our results demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, generalizing across diverse HAR datasets. We believe this work establishes a foundation for future research on time-series and text alignment, paving the way for foundation models in sensor data analysis. Our codes are available at https://github.com/zechenli03/SensorLLM.
[Uncaptioned image] SensorLLM: Aligning Large Language Models with Motion Sensors for Human Activity Recognition
Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2 1University of New South Wales, Sydney 2University of Tokyo {zechen.li, s.deldari, hao.xue1, flora.salim}@unsw.edu.au chen-linyao217@g.ecc.u-tokyo.ac.jp
@@ 1 Introduction
Human Activity Recognition (HAR) is a time-series classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities. Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Wu et al., 2023b; Yin et al., 2023), including text and images.
Enabling LLMs to process sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on time-series data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024). The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—LLM tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024). (ii) Sequence length constraints—sensor data often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs. (iii) Multi-channel complexity—LLMs process univariate inputs, making it difficult to encode multi-sensor data in a way that retains inter-channel dependencies. (iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
To address these challenges, we propose SensorLLM, a framework that aligns sensor data with natural language, allowing LLMs to analyze sensor data through text-based interactions (see Figure 1) without modifying the LLM itself. A major challenge is the annotation bottleneck, further compounded by the complexity and heterogeneity of sensor data. Unlike image-text pairs, sensor data comprises multi-channel numerical signals with diverse characteristics, making direct interpretation and alignment difficult.
Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes. In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from time-series data using statistical analyses and predefined templates. This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics. SensorLLM follows a two-stage framework:
Sensor-Language Alignment Stage.
We automatically generate question-answer pairs to align sensor data with text while preserving temporal features using a pretrained encoder. The resulting embeddings are mapped into a space interpretable by the LLM, mitigating issues associated with text-specific tokenization. Additionally, we introduce special tokens for sensor channels, enabling LLMs to effectively capture multi-channel dependencies.
Task-Aware Tuning Stage.
The aligned embeddings are utilized for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen. This design extends LLMs beyond their original training, addressing concerns raised by Tan et al. (2024) regarding their applicability to time-series data. To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks.
To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks. The key contributions of this work are as follows:
@@
  • We propose a fully automated, human-intuitive approach for aligning time-series data with descriptive text, eliminating the need for manual annotations. Using text similarity metrics, human evaluations, and LLM-based assessments, we demonstrate that SensorLLM effectively captures temporal patterns and channel-specific features, enabling robust multimodal understanding.
@@
  • SensorLLM achieves competitive results across five HAR datasets, matching or surpassing state-of-the-art models. Experiments further validate that modality alignment and task-specific prompts significantly enhance the LLM’s ability to interpret and classify sensor data.
@@
  • We show that SensorLLM maintains strong performance in the Task-Aware Tuning Stage, even when applied to datasets distinct from those used during alignment, highlighting its robustness and generalizability in HAR tasks.
@@ 2 Related Work
In this section, we discuss recent developments in leveraging LLMs for time-series data, specifically focusing on two categories: (1) LLMs for time series as text and (2) Multimodal Large Language Models (MLLMs) for sensor data. A broader overview of other related works, including deep learning approaches to HAR and additional LLM-based forecasting methods, is provided in Appendix A.1.
LLMs for Time Series as Text.
While LLMs excel in processing natural language, applying them directly to time-series data poses unique challenges (Spathis and Kawsar, 2024). Certain methods address this by treating time-series signals as raw text, using the same tokenization as natural language. Notable examples include PromptCast (Xue and Salim, 2023), which transforms numeric inputs into textual prompts for zero-shot forecasting, and LLMTime (Gruver et al., 2024), which encodes time-series as numerical strings for GPT-like models. However, due to the lack of specialized tokenizers for numeric sequences, LLMs may fail to capture crucial temporal dependencies and repetitive patterns (Spathis and Kawsar, 2024). To mitigate these issues, several works employ time-series encoders before mapping the resulting embeddings to language model spaces (Liu et al., 2024a; Zhou et al., 2023c; Xia et al., 2024), thus aligning sensor embeddings with textual embeddings in a contrastive or supervised manner.
MLLMs for Sensor Data.
Extending LLMs to non-textual domains has gained traction, particularly through MLLMs that accept inputs beyond text, such as images or speech. For sensor data, the challenge lies in representing continuous signals effectively. Yoon et al. (2024) propose to ground MLLMs with sensor data via visual prompting. Sensor signals are first visualized as images, guiding the MLLM to analyze the visualized sensor traces alongside task descriptions, which also lower token costs compared to raw-text baselines. Similarly, Moon et al. (2023) introduce IMU2CLIP, which aligns inertial measurement unit streams with text and video in a joint representation space. This approach enables wearable AI applications like motion-based media search and LM-based multimodal reasoning, showcasing how sensor data can be integrated into broader multimodal frameworks.
@@ 3 Methods
In this work, we propose SensorLLM, a framework that aligns sensor data with descriptive text through automatically generated question–answer pairs. Our goal is to develop a multimodal model with reasoning capabilities for analyzing wearable sensor data. As shown in Figure 2, SensorLLM comprises three core components: (1) a pretrained LLM, (2) a pretrained time-series (TS) embedder, and (3) an alignment module MLP, all operating in a two-stage framework.
In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text based on user instructions (Liu et al., 2023a) and questions. In the Task-Aware Tuning Stage, a classifier is added on top of the LLM to perform HAR. Crucially, only the MLP in both stages and the classifier in the Task-Aware Tuning Stage are trainable, while the backbone LLM and TS embedder remain frozen. This design yields a highly efficient and lightweight training process, requiring only 5.67% (535.9M) of the parameters to be trainable in the alignment stage and 0.12% (10.5M) in the tuning stage.
@@ 3.1 Sensor-Text Data Generation
Aligning time-series data with text for LLM-based tasks is challenging due to the lack of rich semantic labels beyond class annotations, making manual annotation impractical (Deldari et al., 2024; Haresamudram et al., 2024). While prior works rely on predefined text prototypes (Sun et al., 2024b; Jin et al., 2024a), we aim for a more human-intuitive representation of sensor data.
We argue that time-series data inherently contains semantic patterns that can be expressed through descriptive text, from simple numerical trends to statistical insights. To achieve this, we automatically generate descriptive text by analyzing observed trends and fluctuations in the data. Using predefined templates, we construct diverse question-answer (QA) pairs that capture trend changes while ensuring accuracy and scalability. These templates (Appendix A.2) are randomly combined to enhance diversity. For example:
@@
where T and S denote specific trends and sensor types, and t corresponds to numerical values.
@@ 3.2 Sensor-Language Alignment
As shown in Figure 2 (a), the Sensor-Language Alignment stage employs a generative model to create multimodal sentences that combine single-channel sensor readings with textual descriptions. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of sensor channels and $T$ is the sequence length. Each channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to preserve its unique characteristics. The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, allowing the model to learn from varying temporal patterns and trend variations. This segmentation strategy ensures that both long-term trends and short-term fluctuations are effectively captured in the generated multimodal sentences.
We use Chronos (Ansari et al., 2024) as the TS embedder, generating segment embeddings $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization. Prior to inputting sensor segments into Chronos (Appendix A.3), we apply instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ to standardize the data. Llama3-8B (Touvron et al., 2023) serves as our LLM backbone.
Alignment Module.
@@
To transform TS embeddings $\hat{x}_{s}$ into text-aligned embeddings $\hat{a}_{s}\in\mathbb{R}^{(l+1)\times D}$ for downstream tasks, we introduce an alignment projection module. This module, implemented as a multi-layer perceptron (MLP), first maps sensor embeddings to an intermediate space of dimension $d_{m}$ and then projects them to the target dimension $D$. Formally,
$\hat{a}_{s}=\mathbf{W}_{2}\cdot\sigma(\mathbf{W}_{1}\hat{x}_{s}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (1)
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{m}\times d_{ts}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{D\times d_{m}}$ are learnable weights, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are biases, and $\sigma$ is the GELU activation function (Hendrycks and Gimpel, 2016). This projection ensures that the transformed embeddings $\hat{a}_{s}$ are semantically aligned with the text embedding space, making them suitable for tasks such as text generation and classification.
Input Embedding.
@@
To integrate sensor data into the LLM, we introduce two special tokens per sensor channel (e.g., <x_acc_start> and <x_acc_end> for the x-axis accelerometer), extending the LLM’s embedding matrix from $\mathbf{E}\in\mathbb{R}^{V\times D}$ to $\mathbf{E}\in\mathbb{R}^{V^{\prime}\times D}$, where $V^{\prime}=V+2c$, with $V$ as the vocabulary size and $c$ as the number of channels. These special token embeddings are concatenated with the aligned sensor embeddings. The final combined sensor representation $\hat{o}_{s}\in\mathbb{R}^{(l+3)\times D}$ is then concatenated with instruction and question embeddings to form the full input sequence $\hat{z}\in\mathbb{R}^{k\times D}$, where $k$ is the total number of tokens.
Loss Function.
@@
SensorLLM processes an input sequence $\mathbf{Z}_{s}=\{z_{s}^{i}\}_{i=1}^{K}$ consisting of sensor and text embeddings and generates an output sequence $\mathbf{Z}_{t}=\{z_{t}^{i}\}_{i=1}^{N}$, where $z_{s}^{i},z_{t}^{i}\in V^{\prime}$, and $K$ and $N$ represent the number of input and output tokens, respectively. The model is trained using a causal language modeling objective, predicting the next token based on previous ones. The optimization minimizes the negative log-likelihood:
$\mathcal{L}_{gen}=-\sum_{i=0}^{N-1}\log P(z_{t}^{i}|Z_{t}^{<i},z_{s}).$ (2)
Loss is computed only on generated tokens, ensuring SensorLLM effectively integrates sensor and text embeddings to produce coherent, contextually appropriate responses.